# KAN-Koopman Based Rapid Detection Of Battery Thermal Anomalies With Diagnostics Guarantees

### v1

@@ 5.1 Details of KAN-Koopman implementation
KAN architecture: We consider a two-layered KAN network with a network configuration of [[4,3,1]][\![4,3,1]\!]: input layer with width 4 (four features $T_{2},T_{\infty},I,\dot{Q}_{c}$), followed by a KAN layer of width 3, and the final KAN layer of width 1 (single output $\mathbb{T}_{1}$). This architecture contains a total of $15$ learnable activation functions, $12$ in the $1^{st}$ and $3$ in the $2^{nd}$ KAN layer. Furthermore, we consider the grid $G=5$ and spline order $\kappa=3$ to approximate the learnable functions.
@@
5.2 Case study I: Incipient thermal fault
@@
5.3 Case study II: Battery charging cyberattack
Figure 2: Under battery charging cyberattack, top plot shows $T_{1}$, $T_{2}$, and $\mathbb{T}_{1}$; second plot shows nominal and corrupted II; third plot shows $\dot{Q}$; fourth and fifth plots, respectively, shows the diagnostic residuals and the thermal anomaly flag generated by proposed and the baseline Koopman algorithm.
For this case study, we consider that the battery is charging at $2C$ rate, and an adversary then injects a $10\%$ bias attack of 0.46A-0.46A from 700s to 1200s. The top plot of Fig 2 shows the true core and surface temperature $T_{1}$, $T_{2}$, and the KAN-estimation $\mathbb{T}_{1}$ under this compromised charging scenario. The second plot of Fig 2 shows the nominal and corrupted charging current, and the third plot captures the additional heat generation under corrupted charging. In this scenario, our diagnostic residual $\mathbb{r}$ crosses the threshold at 708s and thus, reliably detects the presence of thermal anomaly after 8s of attack injection (as shown in the fourth and last plot of Fig 2). The baseline Koopman algorithm detects the thermal anomaly at 720s, leading to a 12s delay in detection compared to the proposed method. Additionally, after the attack withdrawal, residuals for both methods go below the threshold and the flag is reset at 1239s and 1315s as shown the last plot of Fig 2. Thus, the KAN-Koopman method exhibits $60\%$ faster detection and also $64\%$ faster recovery.
@@ 6 Conclusion
In this paper, we propose a model-free KAN-Koopman diagnostic algorithm for the detection of battery thermal anomalies that can be induced from both physical faults and cyberattacks. The proposed method adopts a lightweight KAN network that estimates the core temperature, and next, a Koopman model is learned online using the KAN estimation along with the limited available data to generate predictions for the battery core and surface temperatures. This online learning approach leads to improved adaptability with changes in battery dynamics, while the KAN–Koopman integrated structure ensures rapid detection of battery thermal anomalies. The diagnostic residual is generated as the error in the Koopman predictions for the KAN-based core temperature estimation and surface temperature measurement. Furthermore, we derive analytical conditions for diagnostic guarantees on our KAN-Koopman detection scheme. We compare the performance of our model with the baseline Koopman-only model, that only utilizes surface temperature for diagnostics. Our comparison result shows that the proposed KAN-Koopman method exhibits $24\%$ faster detection of an incipient thermal fault induced thermal anomalies, and $60\%$ faster detection of anomalous heat generation under corrupted charging compared to the Koopman-only model.

### v2

@@ 5.1 Details of KAN-Koopman implementation
KAN architecture: We consider a two-layered KAN network with a network configuration of [[4,3,1]][\![4,3,1]\!]: input layer with width 4 (four features $T_{2},T_{\infty},I,\dot{Q}_{c}$), followed by a KAN layer of width 3, and the final KAN layer of width 1 (single output $\mathbb{T}_{1}$). Furthermore, we consider the grid $G=5$ and spline order $\kappa=3$ for the KAN model.
@@
5.2 Detection of incipient thermal fault
@@ 6 Conclusion
In this paper, we propose a model-free KAN-Koopman diagnostic algorithm for the detection of battery thermal anomalies that can be induced from both physical faults and cyberattacks. The proposed method adopts a lightweight KAN network that estimates the core temperature, and next, a Koopman model is learned online using the KAN estimation along with the limited available data to generate predictions for the battery core and surface temperatures. This online learning approach leads to improved adaptability with changes in battery dynamics, while the KAN–Koopman integrated structure ensures rapid detection of battery thermal anomalies. The diagnostic residual is generated as the error in the Koopman predictions for the KAN-based core temperature estimation and surface temperature measurement. Furthermore, we derive analytical conditions for diagnostic guarantees on our KAN-Koopman detection scheme. We compare the performance of our model with baseline Koopman-only model, that only utilizes surface temperature for diagnostic. Our comparison result shows that the proposed KAN-Koopman method exhibits $24\%$ faster detection of an incipient thermal fault induced thermal anomalies compared to the Koopman-only model.